# Quantumness and Randomness of Quantum Random Number Generators via Photon Statistics

### v2

@@ II Classical vs Quantum Light: Photon-counting Statistics
It is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance $=2$. But, it is not Poisson. This motivates us to propose a two-fold method to test the quantumness of the photon source. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson.
III Testing quantumness by two-fold Statistical Method
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (that are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
@@ III Testing quantumness by two-fold Statistical Method
$=P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}).$
@@
Therefore, $P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})=1-\epsilon$. So, we can say that $(\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}},\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})$ is an approximate confidence interval for the distribution mean $m$ with confidence level $1-\epsilon$.
@@
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s}\geq 1-\frac{y_{\epsilon}}{\sqrt{n}}$, or $\frac{\langle x\rangle}{s}\leq 1+\frac{y_{\epsilon}}{\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Equation (6).
@@
Alternatively, one can perform hypothesis testing to determine whether the light is sub-Poissonian, Poissonia, or super-Poissonian.
Hypothesis testing is a statistical procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample Casella and Berger (2002). A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
@@
To determine the quantumness one has to perform two consecutive hypothesis testing. For both the tests $H_{0}:m=s^{2}$. However, the alternative hypotheses $H_{a}$ would be $m<s^{2}$, and $m>s^{2}$ for the first and second test respectively. For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test, if $z\geq y_{\epsilon}$, where $y_{\epsilon}$ is given by Equation (6), $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian. But if $z<y_{\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform second hypothesis testing with the same statistic. If $z\geq y_{\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian. Note that, this hypothesis testing is equivalent to the above interval estimation technique, and both results in the same decision.
@@
As we saw in Section II, the equality of mean and variance may not imply Poisson distribution. Therefore to ensure the Poisson distribution which is an indicator of a coherent source of light, we suggest a second test which is a goodness-of-fit test, that is, a hypothesis testing to decide whether a sample data comes from a Poisson distribution or not. Chi-square ($\chi^{2}$) testCochran (1952), Kolmogorov–Smirnov test (KS test) Massey (1951), and Anderson–Darling test Anderson and Darling (1952) are some examples of the goodness-of-fit test Dodge (2008a). Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well Conover (1972). The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions Slakter (1965).
@@
Figure 1: Probability density function for $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
@@
H0: The data follow the distribution (15).
[…]
@@
is the well known Gamma function Hogg and Craig (2019). The cumulative distribution function in this case is given by,
@@
Figure 1 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. The area of the critical region is given by $P(\chi^{2}>\chi^{2}_{k,\alpha})=1-F_{\chi^{2}}(\chi^{2}_{k,\alpha})=\alpha$. $\chi^{2}$-test being a one-tailed test, $p$-value for this test is given by $P(\chi^{2}>\chi^{2}_{obs})=1-F_{\chi^{2}}(\chi^{2}_{obs})$, where $\chi^{2}_{obs}$ is the observed value of the test statistic.
@@
If the typical values of the parameters $\mu,T$ in Equation (4) are known, we have $k=N-1$. However, if they are unknown, in that case, $k=N-2$. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if $p$-value is less than $\alpha$.
@@
As mentioned in Section II, in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test for the completeness of the discussion.
IV Testing Quantum Randomness of QRNG: Ideal vs Practical
IV.1 QRNG based on single-photon source
QRNG, based on single photons, uses $/50$ 50 beam splitter Jennewein et al. (2000), where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Figure 2.
[…]
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section III. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with this single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$. So, ideally it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. For example, consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be $\frac{d_{0}}{d_{0}+d_{1}}\frac{T}{100}$. Similarly, the probability of $1$ would be $\frac{d_{1}}{d_{0}+d_{1}}\frac{R}{100}$. If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), ENT Walker (2008) etc. may be used. See Appendix B for a discussion on these.
@@ IV.1 QRNG based on single-photon source
IV.2 QRNG based on Coherent Pulse
Producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999). However, different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023). In this process, the photons follow Poissonian statistics.
A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Figure 3.
SourceAttenuatorSPDFPGARandom Number
Figure 3: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval is negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The FPGA provides random numbers based on the registered time stamps.
If $\lambda$ is the wavelength of the photon wave and $n$ photons are present within $T$ time, the corresponding photon power would be $P=\frac{n\hbar c}{\lambda T}=\frac{\mu\hbar c}{\lambda}$, where $\hbar$ is the Planck’s constant and $c$ is the speed of light in vacuum. Let $P_{a}$ be the photon power coming out from the attenuators. Then, the expected number of photons per unit time is given by $\mu=\frac{P_{a}\lambda}{\hbar c}$. This attenuated pulse is detected by the SPD. The time of this detection is registered to produce photon arrival time-based random numbers.
[…]
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Figure 4. In Figure 4(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows. If $\Delta t_{2i-1}<\Delta t_{2i}$, the $i$-th bit of the random number is taken as 00. If $\Delta t_{2i-1}>\Delta t_{2i}$, the $i$-th bit is taken as $1$. But in this process the random number generation rate is slow. The length of the random bit-string is half of the photon count. Comparatively fast processes have been developed in the literature Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Nie et al. (2014). In Figure 4(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by $P(t_{w}>t)=P(\text{no photon in }(0,t])=P(0,t)=e^{-\mu t}$. Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010). To overcome this drawback, a new scheme is proposed in Nie et al. (2014). In this scheme, as depicted in Figure 4(III), an external time reference is used. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
$=\frac{1}{N^{n}}.$ (10)
IV.3 Time-bin Statistics for Poissonian Photon with Ideal Devices
@@
Currently available photon arrival time-based QRNGs use coherent sources, which generate photons that follow Poissonian statistics. Here we discuss the actual photon statistics in the practical scenario for external reference time-based QRNG (Figure 4(III)) to see the randomness of the generated random number. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@
$=0+e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i-1}{N}\right)\right]^{n}}{n!}$
$\qquad-e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i}{N}\right)\right]^{n}}{n!}$
@@
IV.4 Time-bin Statistics for Poissonian Photon with Erroneous Devices
@@
$P_{d}(i;1st)=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (14)
The remaining probability $1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ is the probability of detecting no photon in a complete reference cycle. In that case, this reference cycle has no contribution to the random number. Therefore the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins. Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detector efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Figure 5 shows how $\mu Td$ affects the probability (15). Here we consider $N=256$. From this plot, it is clear that the randomness increases as $\mu Td$ decreases.
Next, let us consider QRNG in Figure 4(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin, is given by
@@
Since in the probability distribution of Equation (14), $\frac{T}{N}$ is the length of the time-bins, Equation (17) is identical with Equation (14). Therefore, the probability mass function corresponding to the probability Equation (17) is also given by Equation (15) with $T=Nt_{i}$.
@@
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Equation (10)). Since the distributions given by Equation (15) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG.
@@
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution given by Equation (15) is $\epsilon$-random. Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
@@
We call a probability distribution $P(\cdot)$, with support $S$, as $\epsilon$-random if and only if $|P(i)-P(j)|<\epsilon$, for all $i,j\in S$. Since the distribution (15) is a monotone decreasing function, it becomes $\epsilon$-random if and only if
$f(1)-f(N)<\epsilon$
$\implies\frac{e^{\frac{\mu Td}{N}}-1}{e^{\mu Td}-1}\left(e^{\left(1-\frac{1}{N}\right)\mu Td}-1\right)<\epsilon.$ (18)
@@
Figure 6 is a visual representation of the relation (18). It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to generate numbers with $\epsilon$-randomness, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (18) has been satisfied.

### v3

@@ V Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section II.1 we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance $=2$. But, it is not Poisson.
Also, in experimental methods the means and variances are calculated from finite data samples. The fineness of the sample cannot perfectly reflect the mean and variances of the underlying distribution. Therefore, the comparison between the mean and the variance with equality may not reflect the underlying distribution undoubtedly.
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson.
V.1 Interval Estimation
@@
$=P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}).$ (14)
Therefore, combining Eq. (13) and Eq. (14), we have
$P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})=1-\epsilon.$ (15)
@@
So, we can say that $(\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}},\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})$ is an approximate confidence interval for the distribution mean $m$ with confidence level $1-\epsilon$.
@@
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s}\geq 1+\frac{y_{\epsilon}}{\sqrt{n}}$, or $\frac{\langle x\rangle}{s}\leq 1-\frac{y_{\epsilon}}{\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Eq. (13). That is, for some given $\epsilon$ and $y_{\epsilon}$ given by Eq. (13),
[…]
with confidence level $1-\epsilon$.
@@
Alternatively, one can perform hypothesis testing to determine whether the light is sub-Poissonian, Poissonian, or super-Poissonian.
V.2 Hypothesis Testing
To determine the quantumness one has to perform two consecutive hypothesis testing. For the first test,
$H_{0}:\ m=s^{2},\ H_{a}:\ m<s^{2},$ (17)
and for the second test,
$H_{0}:\ m=s^{2},\ H_{a}:\ m>s^{2}.$ (18)
For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test,
case 1: if $z\geq y_{\epsilon}$, where $y_{\epsilon}$ is given by Eq. (13), $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian.
@@ V.2 Hypothesis Testing
case 2: if $z<y_{\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform second hypothesis testing with the same statistic. After the second test, if $z\geq y_{\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian.
Note that, this hypothesis testing is equivalent to the above interval estimation technique, and both results in the same decision.
V.3 Goodness-of-fit Test
@@
Since the equality of mean and variance may not imply Poisson distribution, to ensure the Poisson distribution which is an indicator of a coherent source of light, we suggest a second test which is a goodness-of-fit test, that is, a hypothesis testing to decide whether a sample data comes from a Poisson distribution or not.
@@
Figure 4: Probability density function for $\chi^{2}$ distribution. The critical region, $\chi^{2}>\chi^{2}_{k,\alpha}$, for the $\chi^{2}$ test is denoted by stars, and the region, $\chi^{2}>\chi^{2}_{obs}$, corresponding to the $p$-value is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
@@
$H_{0}$: The data follow the distribution (4).
$H_{\text{a}}$: The data do not follow the distribution (4).
@@
is the well-known Gamma function Hogg and Craig (2019). The cumulative distribution function in this case is given by,
@@
Figure 4 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. The area of the critical region is given by
$P(\chi^{2}>\chi^{2}_{k,\alpha})=1-F_{\chi^{2}}(\chi^{2}_{k,\alpha})=\alpha.$ (22)
$\chi^{2}$-test being a one-tailed test, $p$-value for this test is given by
$P(\chi^{2}>\chi^{2}_{obs})=1-F_{\chi^{2}}(\chi^{2}_{obs}),$ (23)
where $\chi^{2}_{obs}$ is the observed value of the test statistic.
@@
If the typical values of the parameters $\mu,T$ in Eq. (4) are known, we have $k=N-1$. However, if they are unknown, in that case, $k=N-2$. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if $p$-value is less than $\alpha$.
@@
As in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test for the completeness of the discussion.
@@
VI Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (11) of Section IV consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section VII.
VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
@@
Currently available photon arrival time-based QRNGs Qutools (2010); QNu Labs (2018) use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@
$=0+\sum_{n>0}\frac{e^{-\mu T}(\mu T)^{n}}{n!}\left[\left(1-\frac{i-1}{N}\right)^{n}-\left(1-\frac{i}{N}\right)^{n}\right]$
$=e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i-1}{N}\right)\right]^{n}}{n!}-e^{-\mu T}\sum_{n>0}\frac{\left[\mu T\left(1-\frac{i}{N}\right)\right]^{n}}{n!}$
@@
VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
@@
$P_{d}(i;1st)$ $=e^{-(i-1)\frac{\mu Td}{N}}-e^{-i\frac{\mu Td}{N}}$
$=(e^{\frac{\mu Td}{N}}-1)e^{-i\frac{\mu Td}{N}}.$ (27)
The remaining probability
$1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td}$ (28)
is the probability of detecting no photon in a complete reference cycle. In that case, this reference cycle has no contribution to the random number. Therefore the probability mass function of the distribution corresponding to the detection of the first photon in the $i$-th bin is given by
@@
Figure 5: Plots for the probabilities of detecting the first photon in $256$ time-bins. Probabilities for different values of $\mu Td$ as well as uniform probability distribution have been plotted. $\mu$ is the expected number of photons in unit time, $T$ is the time cycle in the external reference time frame, and $d$ is the detection efficiency. If all the optical components are ideal then the distribution should be given by red dashed line.
Figure 5 shows how $\mu Td$ affects the probability (29). Here we consider $N=256$. From this plot, it is clear that the randomness increases as $\mu Td$ decreases. Therefore, the main goal in designing a QRNG is
$\underset{\mu,T,d}{minimize}\ g(\mu,T,d):=\mu Td,$ (30)
$subject\ to\ \mu>0,T>0,d>0.$
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section VII.
Next, let us consider QRNG in Fig. 3(II). Here, the probability, that the waiting time ($t_{w}$) between the arrival of two consecutive photons lies in the $i$-th bin, is given by
@@
Since in the probability distribution of Eq. (VI.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (32) is identical with Eq. (VI.2). Therefore, the probability mass function corresponding to the probability Eq. (32) is also given by Eq. (29) with $T=Nt_{i}$.
@@
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (11)). Since the distributions given by Eq. (29) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG.
@@
Figure 6: Plot representing the relation between $\mu Td$ and $\epsilon$ such that the probability distribution given by Eq. (29) is $\epsilon$-random. Randomness reduces as $\mu Td$ increases. This product can be reduced with the proper choice of external reference cycle $T$, and proper tuning of the source, the attenuators and the detector.
Definition 2.
@@
We define a probability distribution $P(\cdot)$, with support $S$, as $\epsilon$-random if and only if
$\max_{i,j\in S}|P(i)-P(j)|=\epsilon.$ (33)
Since the distribution (29) is a monotone decreasing function, it becomes $\epsilon$-random if and only if
$f(1)-f(N)=\epsilon$
$\implies\frac{e^{\frac{\mu Td}{N}}-1}{e^{\mu Td}-1}\left(e^{\left(1-\frac{1}{N}\right)\mu Td}-1\right)=\epsilon.$ (34)
@@
Figure 6 is a visual representation of the Eq. (34). It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to generate numbers with $\epsilon$-randomness, we need to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (34) has been satisfied.